Vista Electronics, Inc. manufactures two different types of coils used in electric motors. In the fall of the current year, Erica Becker, the controller, compiled the following data.

  • Sales forecast for 20x0 (all units to be shipped in 20x0):
Product Units Price
Light coil 51,000 $ 210
Heavy coil 31,000 310
  • Raw-material prices and inventory levels:
Raw Material Expected Inventories, January 1, 20x0 Desired Inventories, December 31, 20x0 Anticipated Purchase Price
Sheet metal 29,000 lb. 33,000 lb. $ 18
Copper wire 26,000 lb. 29,000 lb. 9
Platform 5,100 units 6,100 units 5
  • Use of raw material:
Amount Used per Unit
Raw Material Light Coil Heavy Coil
Sheet metal 4 lb. 5 lb.
Copper wire 2 lb. 3 lb.
Platform 1 unit
  • Direct-labor requirements and rates:
Product Hours per Unit Rate per Hour
Light coil 4 $ 15
Heavy coil 6 20
  • Finished-goods inventories (in units):
Product Expected January 1, 20x0 Desired December 31, 20x0
Light coil 16,000 21,000
Heavy coil 9,000 10,000
  • Production overhead:
Overhead Cost Item Activity-Based Budget Rate
Purchasing and material handling $ 0.70 per pound of sheet metal and copper wire purchased
Depreciation, utilities, and inspection $ 6.00 per coil produced (either type)
Shipping $ 2.00 per coil shipped (either type)
General production overhead $ 4.00 per direct-labor hour
  1. Prepare the raw material purchases budget (in quantities) for 20x0.
Answer #1

Ans:Raw materials purchases Budget( in quantities)

Particulars Sheet Materials Copper wires Platforms
Materials Required for Light Oil (56,000*4)= $224,000 (56,000*2)= $ 112,000 Nil
Materials required for heavy oil (32,000* 5)=$160,000 (32,000*3)= $96,000 (32,000*1)=$32,000
Total Materials required $ 384,000 $ 208,000 $32,000
Less: Opening inventories (29,000) (26,000) (5,100)
Add: Closing Inventories 33,000 29,000 6,100
Materials required to purchase $388,000 $211,000 $33,000
Total Materials Required to purchase $388,000+$211,000+$33,000= $632,000

Materials Required for Light Coil = 51,000-16,000+21,000 = 56,000

Materials required for Heavy Coil= 31,000- 9,000+10,000 = 32,000